# xTED: Cross-Domain Adaptation via Diffusion-Based Trajectory Editing

### v2

@@ 5.6 Ablation Findings on Model Architecture
Target Data Source Dynamics TC FC Inv CM Ours HC-MR Gravity 23.8$\pm$3.2 25.6$\pm$3.1 28.8$\pm$3.4 20.1$\pm$3.2 31.3$\pm$2.9 Friction 21.9$\pm$3.7 22.4$\pm$4.0 27.9$\pm$3.6 13.1$\pm$3.4 31.8$\pm$3.1 Thigh Size 27.6$\pm$3.5 26.6$\pm$4.1 30.9$\pm$2.7 19.3$\pm$2.1 33.0$\pm$3.0 WK-MR Gravity 23.1$\pm$9.8 15.9$\pm$6.9 20.8$\pm$6.3 11.4$\pm$5.1 23.3$\pm$9.1 Friction 20.4$\pm$6.6 20.7$\pm$8.5 20.7$\pm$7.6 11.4$\pm$5.4 25.9$\pm$9.1 Thigh Size 20.8$\pm$7.5 20.3$\pm$6.2 20.3$\pm$9.0 9.6$\pm$4.0 25.9$\pm$9.1 Total 137.6 131.5 149.4 84.9 171.2
Target Data Source Dynamics TC (Return Cond.) FC (Return Cond.) Ours (Return Cond.) HC-MR Gravity 23.8$\pm$3.2 21.3$\pm$6.2 $\downarrow$ 25.6$\pm$3.1 19.6$\pm$6.5 $\downarrow$ 31.3$\pm$2.9 30.5$\pm$3.1 $\downarrow$ Friction 21.9$\pm$3.7 23.8$\pm$5.6 $\uparrow$ 22.4$\pm$4.0 20.1$\pm$7.9 $\downarrow$ 31.8$\pm$3.1 35.1$\pm$1.8 $\uparrow$ Thigh Size 27.6$\pm$3.5 21.1$\pm$7.7 $\downarrow$ 26.6$\pm$4.1 24.9$\pm$6.4 $\downarrow$ 33.0$\pm$3.0 33.3$\pm$3.0 $\uparrow$ WK-MR Gravity 23.1$\pm$9.8 8.9$\pm$5.2 $\downarrow$ 15.9$\pm$6.9 7.5$\pm$4.2 $\downarrow$ 23.3$\pm$9.1 24.5$\pm$9.0 $\uparrow$ Friction 20.4$\pm$6.6 9.0$\pm$5.5 $\downarrow$ 20.7$\pm$8.5 7.2$\pm$3.8 $\downarrow$ 25.9$\pm$9.1 22.3$\pm$8.7 $\downarrow$ Thigh Size 20.8$\pm$7.5 11.9$\pm$6.3 $\downarrow$ 20.3$\pm$6.2 10.0$\pm$4.8 $\downarrow$ 25.9$\pm$9.1 31.1$\pm$11.5 $\uparrow$ Total 137.6 96.0 $\downarrow$ 131.5 89.3 $\downarrow$ 171.2 176.8 $\uparrow$
We ablate our model architecture to demonstrate the impact of different design insights and techniques. We primarily consider four distinct architecture baselines: (1) Feature Concatenation (FC): concatenating state, action, and reward sequences on feature dimension before passing through attention; (2) Transition concatenation (TC): concatenating state, action, and reward sequences on temporal dimension before passing through attention; (3) Causal Mask (CM): incorporating causal masks in attention modules in our architecture for modeling Markovian process, considering transitions only depending on the previous step; (4) Diffusing over States + Inverse Dynamics (Inv): solely diffusing over state trajectories, while employing two inverse models to obtain actions and rewards. Table 3 demonstrates that our model consistently achieves superior performance, from which several key insights into architecture design can be derived: (1) Entire trajectory modeling outperforms partial trajectory modeling: Using inverse models (Inv) results in significantly higher compounding errors, as action feasibility and reward plausibility cannot be reliably maintained when only states are diffused without a unified model that handles all components consistently; (2) Imposing Markovian properties as priors in trajectory modeling is suboptimal: Incorporating Markovian modeling (CM) degrades the quality of edited source data, as local transition dynamics enforced by masked attention mechanisms may conflict with the global temporal consistency required in trajectory modeling. This suggests that trajectory modeling demands a distinct approach compared to common practices in decision-making approaches; (3) Our design effectively captures intricate dependencies within trajectories: Concatenating states, actions, and rewards along transition or feature dimensions (TC/FC) is problematic. It treats the entire trajectory as a homogeneous, image-like input, which obscures the model’s ability to correctly identify dependencies between transitions and decision elements, and introduces unwarranted dependencies of rewards on state-action sequences; (4) Our design integrates well with return-conditioned diffusion: Table 4 reveals that incorporating return conditioning in naïve architectures (TC/FC) leads to performance degradation, whereas our model with return conditioning successfully guides edited source data towards higher-reward regions, resulting in superior policy performance. This is because only our design correctly models the one-way dependencies of reward and state-action sequences, enabling effectively implementing return-conditioned strategies.
@@ A.1 Domain Adaptation
Existing categories of domain adaptation.
The first category of approaches to achieving such alignment is through learning direct mappings or corrections between source and target domains. For observation space alignment, numerous works leverage advances in generative adversarial learning and contrastive learning to transform the observation space, achieving consistent visual rendering and camera viewpoints as in the target domain (Zhu et al., 2017; Liu et al., 2018; Zhang et al., 2019; Bewley et al., 2019; Rao et al., 2020; Liu et al., 2020; Gangwani & Peng, 2020; Radosavovic et al., 2021). In scenarios with dynamics mismatch, it becomes necessary to align both state and action spaces simultaneously, which introduces more complex correspondence learning mechanisms such as (dynamics/temporal) cycle consistency (Kim et al., 2020; Zhang et al., 2020; Raychaudhuri et al., 2021; Zakka et al., 2021; Wang et al., 2022b) and mutual information criteria (Franzmeyer et al., 2022). Alternatively, modifying the reward space can compensate for the misalignment in transitional dynamics within MDPs (Eysenbach et al., 2021; Liu et al., 2022; Lyu et al., 2024; Xue et al., 2023), which may also be implicitly addressed by adaptively regularizing Q-values (Niu et al., 2022; 2023; Hou et al., 2024). This category necessitates model retraining when new data from other source domains becomes available, which can be time-consuming and labor-intensive.
The second category resorts to building a canonical representation space that preserves task-relevant yet domain-agnostic information (Stadie et al., 2016; Pan et al., 2017; Mueller et al., 2018; Sermanet et al., 2018; Cetin & Celiktutan, 2021; Wang et al., 2022a; Yang et al., 2023; Choi et al., 2024). This auxiliary representation space, often viewed as skills (Gupta et al., 2017; Hejna et al., 2020; Pertsch et al., 2022; Xu et al., 2023b) or subgoals (Sharma et al., 2019; Ma et al., 2022; 2023; Li et al., 2024), facilitates the construction of aligned domain-invariant MDPs by capturing task semantics that effectively guide agents in completing tasks. However, this approach involves designing task-specific loss functions and optimization procedures, which may struggle to find a representation space that effectively encapsulates the inputs and reflects the task.
@@ A.2 Distinctions from Data Augmentation
We have to emphasize that using diffusion models for data editing is both sample-efficient and error-tolerant than diffusion-based trajectory generation for data augmentation (Lu et al., 2023; He et al., 2023; Jackson et al., 2024). This is because the goal in data editing is far more easily attainable compared to its application in data generation tasks. When diffusion models are used for data generation, the focus is usually on achieving a high level of precision to generate data that closely matches the original data distribution. However, in the context of data editing, the requirements are relaxed to simply distort (edit) the data to more closely match the target dynamics and observation patterns than the source data. Crucially, there is no need for edited source data to closely resemble target data in every detail. Thus, diffusion models for editing are relatively insensitive to the amount of target data used, having minimal impact on the ultimate policy performance as proved in Section 5.5.
@@ B.1 Simulation Experiments
    The computation of attention modules is proportional to $9H^{2}$ (self-attention: $(N_{s}+N_{a}+N_{r})^{2}$; cross-attention:$4H^{2}$), so we also set architecture with minimal differences on computation efforts for the architecture ablation baselines for fair comparisons, i.e. FC, TC, Inv, CM. For CM, we keep the architecture the same with xTED, with the only difference of adding a causal mask with markovian constraints. We use 9 successive self-attention modules and FC ($9\times H^{2}$), while one self-attention module for TC ($(3\times H)^{2}$), achieving the same attention computation. For Inv, we only have to diffuse over states without addressing the intricate dependencies within trajectories so we reduce the computation and adopt only 2 self-attention modules for efficiency.
@@ B.2 Real-Robot Experiments
  • Model and Method Setups: Different from simulation experiments, the robot manipulation tasks involves environments where rewards are hard to specify, thus resulting in collecting non-reward trajectories. Correspondingly, the self-attention and cross-attention modules for reward trajectories are removed from the original xTED architecture, leaving only self-attention and cross-attention modules for state and action sequences. In terms of other architecture designs and editing details, we keep them identical with simulation experiment setups. We set the trajectory horizon for modeling with xTED to $H=5$, as the robot manipulation trajectories contain higher-dimensional features, necessitating a trade-off in horizon length to reduce computational burden.
@@ C.1 Numeric Results of Real-Robot Experiments
Target Data Source Dynamics $e=1,\kappa=1.0$ $e=1,\kappa=0.5$ $e=2,\kappa=0.5$ WK-MR Gravity 27.1$\pm$10.8 23.3$\pm$9.1 28.3$\pm$9.2 Friction 25.5$\pm$13.6 25.9$\pm$9.1 26.8$\pm$8.9 Thigh Size 24.1$\pm$9.3 25.9$\pm$9.1 24.7$\pm$11.0
@@
Target Data Source Dynamics Target Target+Source $\Delta$ Target+Edited Source $\Delta$ HC-MR Gravity 26.2$\pm$3.5 30.9$\pm$2.9 +17.9% 33.9$\pm$2.3 +29.4% Friction 26.2$\pm$3.5 26.4$\pm$3.3 +0.8% 28.0$\pm$3.2 +6.9% Thigh Size 26.2$\pm$3.5 23.7$\pm$5.4 -9.5% 31.0$\pm$3.7 +18.3% WK-MR Gravity 16.6$\pm$5.9 24.5$\pm$13.6 +47.6% 26.7$\pm$14.8 +60.8% Friction 16.6$\pm$5.9 27.9$\pm$10.8 +68.1% 36.2$\pm$11.7 +118.1% Thigh Size 16.6$\pm$5.9 21.7$\pm$12.1 +30.7% 27.6$\pm$14.7 +66.3% Total 128.4 155.1 +20.8% 183.4 +46.6%
@@ C.5 Comparisons on Transformer and Temporal U-Net for Diffusion-Based Data Editing
Target Data Source Dynamics Temporal U-Net xTED
@@
We train xTED with constraints of fixed initial transition (Init. Const.) and no constraint of fixed last transition (Last Const.) in trajectories. xTED can also accommodate conditioning strategies (Conditioned xTED). In this section, we ablate on these additional constraints and conditions.
@@
Target Data Source Dynamics w/o Init. Const. w/ Last Const. xTED Conditioned xTED HC-MR Gravity 28.6$\pm$4.3 21.8$\pm$3.3 31.3$\pm$2.9 30.5$\pm$3.1 Friction 26.3$\pm$4.1 22.0$\pm$3.7 31.8$\pm$3.1 35.1$\pm$1.8 Thigh Size 29.6$\pm$3.4 21.3$\pm$3.9 33.0$\pm$3.0 33.3$\pm$3.0 WK-MR Gravity 24.5$\pm$7.7 18.7$\pm$6.9 23.3$\pm$9.1 24.5$\pm$9.0 Friction 20.6$\pm$7.5 15.2$\pm$6.7 25.9$\pm$9.1 22.3$\pm$8.7 Thigh Size 18.2$\pm$7.3 14.5$\pm$7.0 25.9$\pm$9.1 31.1$\pm$11.5 Total 147.8 113.5 171.2 176.8
@@
Target Data Source Dynamics Target+Edited Source (horizon=20) Target+Edited Source (horizon=10) Halfcheetah Med Gravity 40.6$\pm$2.0 40.1$\pm$2.4 Friction 41.2$\pm$1.6 40.6$\pm$1.6 Thigh Size 40.7$\pm$2.4 40.7$\pm$2.2 Med-R Gravity 31.3$\pm$2.9 28.8$\pm$3.3 Friction 31.8$\pm$3.1 29.8$\pm$2.8 Thigh Size 33.0$\pm$3.0 30.2$\pm$3.2 Med-E Gravity 43.8$\pm$3.6 43.6$\pm$3.6 Friction 43.2$\pm$3.0 37.6$\pm$3.8 Thigh Size 43.0$\pm$3.0 40.6$\pm$3.5 Walker2d Med Gravity 58.2$\pm$11.7 53.1$\pm$12.6 Friction 54.5$\pm$13.7 58.9$\pm$12.5 Thigh Size 58.9$\pm$11.7 57.8$\pm$12.0 Med-R Gravity 23.3$\pm$9.1 21.8$\pm$7.1 Friction 25.9$\pm$9.1 25.0$\pm$9.1 Thigh Size 25.9$\pm$9.1 18.9$\pm$7.5 Med-E Gravity 82.9$\pm$18.1 76.0$\pm$20.0 Friction 74.0$\pm$24.5 73.2$\pm$16.2 Thigh Size 81.0$\pm$21.4 75.1$\pm$22.9 Total 833.2 791.8

### v3

@@ 5.2 Baselines
Target Data Domain Gap Src Tgt DARA DARA+xTED Halfcheetah Med Gravity 8.4$\pm$6.3 39.5$\pm$2.4 40.6$\pm$1.9 41.1$\pm$2.0 Friction 14.4$\pm$3.6 39.5$\pm$2.4 30.8$\pm$6.3 41.4$\pm$1.5 Thigh Size -0.1$\pm$1.0 39.5$\pm$2.4 40.4$\pm$2.4 41.5$\pm$1.6 Med-R Gravity 8.4$\pm$6.3 26.2$\pm$3.5 27.2$\pm$3.2 30.7$\pm$2.9 Friction 14.4$\pm$3.6 26.2$\pm$3.5 23.2$\pm$3.1 31.2$\pm$2.8 Thigh Size -0.1$\pm$1.0 26.2$\pm$3.5 29.1$\pm$4.3 32.9$\pm$2.7 Med-E Gravity 8.4$\pm$6.3 40.1$\pm$2.9 39.7$\pm$3.8 45.3$\pm$3.8 Friction 14.4$\pm$3.6 40.1$\pm$2.9 23.8$\pm$9.5 42.8$\pm$2.4 Thigh Size -0.1$\pm$1.0 40.1$\pm$2.9 37.9$\pm$4.9 41.8$\pm$3.3 Walker2d Med Gravity 16.6$\pm$6.2 45.3$\pm$15.9 53.2$\pm$11.1 59.0$\pm$10.5 Friction 8.2$\pm$1.4 45.3$\pm$15.9 51.1$\pm$11.4 54.3$\pm$10.8 Thigh Size 7.6$\pm$3.5 45.3$\pm$15.9 49.1$\pm$10.2 56.8$\pm$17.6 Med-R Gravity 16.6$\pm$6.2 16.6$\pm$5.9 13.2$\pm$5.7 23.1$\pm$7.1 Friction 8.2$\pm$1.4 16.6$\pm$5.9 28.3$\pm$9.0 25.8$\pm$7.4 Thigh Size 7.6$\pm$3.5 16.6$\pm$5.9 21.8$\pm$7.1 22.1$\pm$9.5 Med-E Gravity 16.6$\pm$6.2 71.0$\pm$21.0 73.4$\pm$14.3 82.5$\pm$18.7 Friction 8.2$\pm$1.4 71.0$\pm$21.0 78.8$\pm$20.2 78.1$\pm$18.3 Thigh Size 7.6$\pm$3.5 71.0$\pm$21.0 83.2$\pm$16.4 83.2$\pm$16.4 Total 165.1 716.1 744.8 833.6
@@
Target Data Domain Gap TC FC Inv CM Ours HC-MR Gravity 23.8$\pm$3.2 25.6$\pm$3.1 28.8$\pm$3.4 20.1$\pm$3.2 31.3$\pm$2.9 Friction 21.9$\pm$3.7 22.4$\pm$4.0 27.9$\pm$3.6 13.1$\pm$3.4 31.8$\pm$3.1 Thigh Size 27.6$\pm$3.5 26.6$\pm$4.1 30.9$\pm$2.7 19.3$\pm$2.1 33.0$\pm$3.0 WK-MR Gravity 23.1$\pm$9.8 15.9$\pm$6.9 20.8$\pm$6.3 11.4$\pm$5.1 23.3$\pm$9.1 Friction 20.4$\pm$6.6 20.7$\pm$8.5 20.7$\pm$7.6 11.4$\pm$5.4 25.9$\pm$9.1 Thigh Size 20.8$\pm$7.5 20.3$\pm$6.2 20.3$\pm$9.0 9.6$\pm$4.0 25.9$\pm$9.1 Total 137.6 131.5 149.4 84.9 171.2
5.5 Ablations on Model Architecture
We ablate our model architecture to demonstrate the impact of different design insights and techniques. We primarily consider four distinct architecture baselines: (1) Feature Concatenation (FC): concatenating state, action, and reward sequences on feature dimension before passing through attention; (2) Transition concatenation (TC): concatenating state, action, and reward sequences on temporal dimension before passing through attention; (3) Causal Mask (CM): incorporating causal masks in attention blocks in our architecture for modeling Markovian process, considering transitions only depending on the previous step; (4) Diffusing over States + Inverse Dynamics (Inv): solely diffusing over state trajectories, while employing two inverse models to obtain actions and rewards.
Table 3 demonstrates that our model consistently achieves superior performance, from which several key insights into architecture design can be derived: (1) Entire trajectory modeling outperforms partial trajectory modeling: Using inverse models (Inv) results in significantly higher compounding errors, as action feasibility and reward plausibility cannot be reliably maintained when only states are diffused without a unified model that handles all components consistently; (2) Imposing Markovian properties as priors in trajectory modeling is suboptimal: Incorporating Markovian modeling (CM) degrades the quality of edited source data, as local transition dynamics enforced by masked attention mechanisms may conflict with the global temporal consistency required in trajectory modeling. This suggests that trajectory modeling demands a distinct approach compared to common practices in decision-making approaches; (3) Our design effectively captures intricate dependencies within trajectories: Concatenating states, actions, and rewards along transition or feature dimensions (TC/FC) is problematic. It treats the entire trajectory as a homogeneous, image-like input, which obscures the model’s ability to correctly identify dependencies between transitions and decision elements, and introduces unwarranted dependencies of rewards on state-action sequences.
@@ A.1 Domain Adaptation
Existing research directions in domain adaptation.
The first category of approaches to achieving such alignment is through learning direct mappings or corrections between source and target domains. For observation space alignment, numerous works leverage advances in generative adversarial learning and contrastive learning to transform the observation space, achieving consistent visual rendering and camera viewpoints as in the target domain (Zhu et al., 2017; Liu et al., 2018; Zhang et al., 2019; Bewley et al., 2019; Rao et al., 2020; Liu et al., 2020; Gangwani & Peng, 2020; Radosavovic et al., 2021). In scenarios with dynamics mismatch, it becomes necessary to align both state and action spaces simultaneously, which introduces more complex correspondence learning mechanisms such as (dynamics/temporal) cycle consistency (Kim et al., 2020; Zhang et al., 2020; Raychaudhuri et al., 2021; Zakka et al., 2021; Wang et al., 2022b) and mutual information criteria (Franzmeyer et al., 2022). Alternatively, modifying the reward space can compensate for the misalignment in transitional dynamics within MDPs (Eysenbach et al., 2021; Liu et al., 2022; Lyu et al., 2024; Xue et al., 2023), which may also be implicitly addressed by adaptively regularizing Q-values (Niu et al., 2022, 2023; Hou et al., 2024). This category necessitates model retraining when new data from other source domains becomes available, which can be time-consuming and labor-intensive.
The second category resorts to building a canonical representation space that preserves task-relevant yet domain-agnostic information (Stadie et al., 2016; Pan et al., 2017; Mueller et al., 2018; Sermanet et al., 2018; Cetin & Celiktutan, 2021; Wang et al., 2022a; Yang et al., 2023; Choi et al., 2024). This auxiliary representation space, often viewed as skills (Gupta et al., 2017; Hejna et al., 2020; Pertsch et al., 2022; Xu et al., 2023b) or subgoals (Sharma et al., 2019; Ma et al., 2022, 2023; Li et al., 2024), facilitates the construction of aligned domain-invariant MDPs by capturing task semantics that effectively guide agents in completing tasks. However, this approach involves designing task-specific loss functions and optimization procedures, which may struggle to find a representation space that effectively encapsulates the inputs and reflects the task.
@@ A.2 Distinctions from Data Augmentation
We have to emphasize that using diffusion models for data editing is both sample-efficient and error-tolerant than diffusion-based trajectory generation for data augmentation (Lu et al., 2023; He et al., 2023; Jackson et al., 2024). This is because the goal in data editing is far more easily attainable compared to its application in data generation tasks. When diffusion models are used for data generation, the focus is usually on achieving a high level of precision to generate data that closely matches the original data distribution. However, in the context of data editing, the requirements are relaxed to simply distort (edit) the data to more closely match the target dynamics and observation patterns than the source data. Crucially, there is no need for edited source data to closely resemble target data in every detail. Thus, diffusion models for editing are relatively insensitive to the amount of target data used, having minimal impact on the ultimate policy performance.
@@ B.1 Simulation Experiments
    The computation of attention blocks is proportional to $9H^{2}$ (self-attention: $(N_{s}+N_{a}+N_{r})^{2}$; cross-attention:$4H^{2}$), so we also set architecture with minimal differences on computation efforts for the architecture ablation baselines for fair comparisons, i.e. FC, TC, Inv, CM. For CM, we keep the architecture the same with xTED, with the only difference of adding a causal mask with markovian constraints. We use 9 successive self-attention blocks and FC ($9\times H^{2}$), while one self-attention block for TC ($(3\times H)^{2}$), achieving the same attention computation. For Inv, we only have to diffuse over states without addressing the intricate dependencies within trajectories so we reduce the computation and adopt only 2 self-attention blocks for efficiency.
@@ B.2 Real-Robot Experiments
  • Model and Method Setups: Different from simulation experiments, the robot manipulation tasks involves environments where rewards are hard to specify, thus resulting in collecting non-reward trajectories. Correspondingly, the self-attention and cross-attention blocks for reward trajectories are removed from the original xTED architecture, leaving only self-attention and cross-attention blocks for state and action sequences. In terms of other architecture designs and editing details, we keep them identical with simulation experiment setups. We set the trajectory horizon for modeling with xTED to $H=5$, as the robot manipulation trajectories contain higher-dimensional features, necessitating a trade-off in horizon length to reduce computational burden.
@@ C.1 Numeric Results of Real-Robot Experiments
Target Data Domain Gap $e=1,\kappa=1.0$ $e=1,\kappa=0.5$ $e=2,\kappa=0.5$ WK-MR Gravity 27.1$\pm$10.8 23.3$\pm$9.1 28.3$\pm$9.2 Friction 25.5$\pm$13.6 25.9$\pm$9.1 26.8$\pm$8.9 Thigh Size 24.1$\pm$9.3 25.9$\pm$9.1 24.7$\pm$11.0
@@
Target Data Domain Gap Target Target+Source $\Delta$ Target+Edited Source $\Delta$ HC-MR Gravity 26.2$\pm$3.5 30.9$\pm$2.9 +17.9% 33.9$\pm$2.3 +29.4% Friction 26.2$\pm$3.5 26.4$\pm$3.3 +0.8% 28.0$\pm$3.2 +6.9% Thigh Size 26.2$\pm$3.5 23.7$\pm$5.4 -9.5% 31.0$\pm$3.7 +18.3% WK-MR Gravity 16.6$\pm$5.9 24.5$\pm$13.6 +47.6% 26.7$\pm$14.8 +60.8% Friction 16.6$\pm$5.9 27.9$\pm$10.8 +68.1% 36.2$\pm$11.7 +118.1% Thigh Size 16.6$\pm$5.9 21.7$\pm$12.1 +30.7% 27.6$\pm$14.7 +66.3% Total 128.4 155.1 +20.8% 183.4 +46.6%
@@ C.5 Comparisons on Transformer and Temporal U-Net for Diffusion-Based Data Editing
Target Data Domain Gap Temporal U-Net xTED
@@
We train xTED with constraints of fixed initial transition (Init. Const.) and no constraint of fixed last transition (Last Const.) in trajectories, like image inpainting techniques (Rombach et al., 2022) in previous diffusion-based image synthesis works that fix some parts of the original image during entire editing/generation process. xTED can also accommodate conditioning strategies (Conditioned xTED). In this section, we ablate on these additional constraints and conditions.
@@
Target Data Domain Gap w/o Init. Const. w/ Last Const. xTED Conditioned xTED HC-MR Gravity 28.6$\pm$4.3 21.8$\pm$3.3 31.3$\pm$2.9 30.5$\pm$3.1 Friction 26.3$\pm$4.1 22.0$\pm$3.7 31.8$\pm$3.1 35.1$\pm$1.8 Thigh Size 29.6$\pm$3.4 21.3$\pm$3.9 33.0$\pm$3.0 33.3$\pm$3.0 WK-MR Gravity 24.5$\pm$7.7 18.7$\pm$6.9 23.3$\pm$9.1 24.5$\pm$9.0 Friction 20.6$\pm$7.5 15.2$\pm$6.7 25.9$\pm$9.1 22.3$\pm$8.7 Thigh Size 18.2$\pm$7.3 14.5$\pm$7.0 25.9$\pm$9.1 31.1$\pm$11.5 Total 147.8 113.5 171.2 176.8
Compared with other model architectures, we find that our design integrates well with return-conditioned diffusion: Table 12 reveals that incorporating return conditioning in naïve architectures (TC/FC) leads to performance degradation, while in our model successfully guides edited source data towards higher-reward regions, resulting in superior policy performance. This is because only our design correctly models the one-way dependencies of reward and state-action sequences, enabling effectively implementing return-conditioned strategies.
Target Data Domain Gap TC (Return Cond.) FC (Return Cond.) Ours (Return Cond.) HC-MR Gravity 23.8$\pm$3.2 21.3$\pm$6.2 $\downarrow$ 25.6$\pm$3.1 19.6$\pm$6.5 $\downarrow$ 31.3$\pm$2.9 30.5$\pm$3.1 $\downarrow$ Friction 21.9$\pm$3.7 23.8$\pm$5.6 $\uparrow$ 22.4$\pm$4.0 20.1$\pm$7.9 $\downarrow$ 31.8$\pm$3.1 35.1$\pm$1.8 $\uparrow$ Thigh Size 27.6$\pm$3.5 21.1$\pm$7.7 $\downarrow$ 26.6$\pm$4.1 24.9$\pm$6.4 $\downarrow$ 33.0$\pm$3.0 33.3$\pm$3.0 $\uparrow$ WK-MR Gravity 23.1$\pm$9.8 8.9$\pm$5.2 $\downarrow$ 15.9$\pm$6.9 7.5$\pm$4.2 $\downarrow$ 23.3$\pm$9.1 24.5$\pm$9.0 $\uparrow$ Friction 20.4$\pm$6.6 9.0$\pm$5.5 $\downarrow$ 20.7$\pm$8.5 7.2$\pm$3.8 $\downarrow$ 25.9$\pm$9.1 22.3$\pm$8.7 $\downarrow$ Thigh Size 20.8$\pm$7.5 11.9$\pm$6.3 $\downarrow$ 20.3$\pm$6.2 10.0$\pm$4.8 $\downarrow$ 25.9$\pm$9.1 31.1$\pm$11.5 $\uparrow$ Total 137.6 96.0 $\downarrow$ 131.5 89.3 $\downarrow$ 171.2 176.8 $\uparrow$
@@
Target Data Domain Gap Target+Edited Source (horizon=20) Target+Edited Source (horizon=10) Halfcheetah Med Gravity 40.6$\pm$2.0 40.1$\pm$2.4 Friction 41.2$\pm$1.6 40.6$\pm$1.6 Thigh Size 40.7$\pm$2.4 40.7$\pm$2.2 Med-R Gravity 31.3$\pm$2.9 28.8$\pm$3.3 Friction 31.8$\pm$3.1 29.8$\pm$2.8 Thigh Size 33.0$\pm$3.0 30.2$\pm$3.2 Med-E Gravity 43.8$\pm$3.6 43.6$\pm$3.6 Friction 43.2$\pm$3.0 37.6$\pm$3.8 Thigh Size 43.0$\pm$3.0 40.6$\pm$3.5 Walker2d Med Gravity 58.2$\pm$11.7 53.1$\pm$12.6 Friction 54.5$\pm$13.7 58.9$\pm$12.5 Thigh Size 58.9$\pm$11.7 57.8$\pm$12.0 Med-R Gravity 23.3$\pm$9.1 21.8$\pm$7.1 Friction 25.9$\pm$9.1 25.0$\pm$9.1 Thigh Size 25.9$\pm$9.1 18.9$\pm$7.5 Med-E Gravity 82.9$\pm$18.1 76.0$\pm$20.0 Friction 74.0$\pm$24.5 73.2$\pm$16.2 Thigh Size 81.0$\pm$21.4 75.1$\pm$22.9 Total 833.2 791.8